OpenStudy (anonymous):

whats the correct simplification of this ? x^9y14/x^2y^9 ?

OpenStudy (anonymous):

use the property of \[\frac{1}{x}=x^{-1}\]

OpenStudy (anonymous):

using this you get \[x^9y^14x^{-2}y^{-9}\]

OpenStudy (anonymous):

then you can use the property \[a^n*a^m=a^{n+m}\]

OpenStudy (anonymous):

really the first property is \[\frac{1}{a^n}=a^{(-1)n}\]

OpenStudy (anonymous):

These are the choices a) x^11 y^11 b) x^12 y^12 c) x^7 y^5 d) x18 y16

OpenStudy (anonymous):

use the properties i showed you if the bases of the exponents are the same you add the exponents

OpenStudy (anonymous):

\[x^9y^{14}x^{-2}y^{-9}=x^{9+-2}y^{14+-9}\]